OpenStudy (anonymous):

Divide. x^2+2x+1/x-2/x^2-1/x^2-4

OpenStudy (anonymous):

[x^2+2x+1/x-2] * [x^2-4/x^2-1] \[\frac{ (x+1)(x+1) }{ x-2 } * \frac{ (x+2)(x-2) }{ (x+1)(x-1) }\] \[\frac{ (x+1)1 }{ 1} * \frac{ (x+2)1 }{ 1(x-1) }\] \[\frac{ (x+1)(x+2) }{ (x-1)}\]

OpenStudy (anonymous):

ty